to save $600 every mommy ina retrement ann ity that much will he be able to withdraw (approximately each month during his 25 year retrement s paying 6.5% til his retirement in 30 years. f the annuity keeps pa ng the same interest 2232.50 4455.75 2.78565
Answer #1

Please find below the answer and please do provide rating.. Let me know if you need any clarification ...

We have compute first Future Value of annuity after 30 year.
Put in financial calculator
PMT -600
I =6.5%/12 =0.54% 0.54%
N =30*12 =360 360
PV 0
Compute FV $661,133.01
We have to compute now the annuity withdrawn for next 25 year every month.
Put in financial calculator below
N=25*12 300
I=6.5%/12 0.54%
FV 0
PV= ($661,133.01)
PMT $4,455.75
There fore John Snow can with draw $ 4455.75 per month for next 25 year
Correct answer is option $4,455.75